OpenStudy (anonymous):

Find the Domain and Range of the following sets of ordered pair: {(x,y) | y=x^2-2} {(x,y) | y=√ 4-x}

OpenStudy (lgbasallote):

okay...so like i said i'll teach you the 2nd restriction

OpenStudy (lgbasallote):

restriction 2 is in form \[y = \sqrt{f(x)}\]

OpenStudy (anonymous):

You always get [Math Processing Error] ?

OpenStudy (lgbasallote):

let's take d) for example..it has a square root right?

OpenStudy (anonymous):

Yep

OpenStudy (lgbasallote):

and is there a variable in the square root?

OpenStudy (anonymous):

there i

OpenStudy (lgbasallote):

therefore it qualifies for restriction 2

OpenStudy (anonymous):

uhuh

OpenStudy (lgbasallote):

for restriction two you will equate the radicand (expression inside the radical) as greater than or equal to 0

OpenStudy (anonymous):

*math nosebleed*

OpenStudy (lgbasallote):

for d) \[\sqrt{4-x}\] you simply do \[4 - x \ge 0\] that's all it means

OpenStudy (lgbasallote):

solve for x

OpenStudy (anonymous):

Uhmmm. You always get this lgbasallote Medals 0 for d) [Math Processing Error] you simply do [Math Processing Error] that's all it means

OpenStudy (anonymous):

?

OpenStudy (lgbasallote):

refresh

OpenStudy (lgbasallote):

it's just an error in the reading of the code i typed which can be easily remedied by refreshing the page

OpenStudy (lgbasallote):

just tell me when you're okay

OpenStudy (anonymous):

Got it. So X would be \[x \ge -4\]

OpenStudy (anonymous):

?

OpenStudy (lgbasallote):

4 - x \(\ge\) 0 my way is to add x to both sides 4 \(\ge\) x therefore \(x \le 4\)the domain is x\(\le|) 4 or all values of x less than or equal to 4

OpenStudy (lgbasallote):

the domain is \(x \le 4\)*

OpenStudy (anonymous):

Okay. NOw i get it. So how about the one with an exponent? C

OpenStudy (anonymous):

Wait. Lol the Range first. for D

OpenStudy (lgbasallote):

lol the one with the exponent doesnt matter...if you look at that equation it looks nothing like the erstrictions so the domain is just all real values (which means no restrictions) icannot say the same for its range though coz i cannot compute it mentally ANYWAY range of d...square both sides to get rid of square root then solve for x

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

\[y ^{2} = 4-x\]

OpenStudy (anonymous):

X = Y^2 +4 ?

OpenStudy (lgbasallote):

good...solve for x

OpenStudy (anonymous):

Do I equate i to 0 AGain?

OpenStudy (anonymous):

X= -4?

OpenStudy (lgbasallote):

oh...no..x = y^2 + 4 is fine...my message was late sent

OpenStudy (lgbasallote):

okay look at y^2 + 4 is it a fraction?

OpenStudy (anonymous):

NOpe

OpenStudy (lgbasallote):

does it have a square root?

OpenStudy (anonymous):

Nope. but it's squared

OpenStudy (lgbasallote):

nahh we're just concerned IF it is a square root

OpenStudy (lgbasallote):

none of the restrictions are satisfied therefore no restirctions...the range is all real numbers

OpenStudy (anonymous):

Phew! That was brain wrecking. Okay

OpenStudy (lgbasallote):

ikr

OpenStudy (lgbasallote):

as for c...i have taught you everything..can you do it yourself?

OpenStudy (lgbasallote):

you can ask me if you have doubts on the steps/answer though

OpenStudy (anonymous):

Uhmm . The domain is a set of real numbers right?

OpenStudy (anonymous):

I still don't get it. I'm a tad confused atm.

OpenStudy (lgbasallote):

yep

OpenStudy (anonymous):

how do you find the range?

OpenStudy (anonymous):

Is it also set of real numbers?

OpenStudy (anonymous):

So when it comes to graphing Function what does it mean if a point it circled?

OpenStudy (lgbasallote):

forst add 2 to both sides so x^2 will be alone

OpenStudy (anonymous):

and then square root both side to get rid of ^2?

OpenStudy (lgbasallote):

yep

OpenStudy (lgbasallote):

so it's restriction 2

OpenStudy (anonymous):

So it's under restriction

OpenStudy (lgbasallote):

yes

OpenStudy (anonymous):

Then?

OpenStudy (anonymous):

I eqaute the radiant to 0?

OpenStudy (anonymous):

for restriction two you will equate the radicand (expression inside the radical) as greater than or equal to 0 ?

OpenStudy (lgbasallote):

yep

OpenStudy (anonymous):

wait. What happens to Y?

OpenStudy (lgbasallote):

what do you get as the expression inside the square root?

OpenStudy (anonymous):

\[x = \sqrt{2y}\]

OpenStudy (lgbasallote):

hmm no...it should be \[\sqrt{y+2}\] recheck your solution and see where you went wrong

OpenStudy (anonymous):

ohh sorry typo

OpenStudy (anonymous):

So it plus. So x = \[\sqrt{y+2}\]

OpenStudy (anonymous):

So do I equate y+ 2 to 0?

OpenStudy (lgbasallote):

y+2 \(\ge\) 0

OpenStudy (anonymous):

Ohhhh. So that' how it goes but how will you know if it's greater than or less than?

OpenStudy (lgbasallote):

it's always greater than or equal to 0 i mean at first

OpenStudy (lgbasallote):

if y or x becomes negative then you switch the sign

OpenStudy (anonymous):

Ohhh. Okay now I get it. Thank!

OpenStudy (anonymous):

Now on for the next one. Do you know I have 12 more question to go. *sighs* and it's all due on june 8. Guess I'll be pulling an all nighter.

OpenStudy (lgbasallote):

well im not o.O sorry lol..pretty sure lots of others here can answer them though

OpenStudy (anonymous):

Yep. thanks for you help! and the medal ^^

OpenStudy (lgbasallote):

you were a good student :p

OpenStudy (anonymous):

:))) You're a good teacher xP